OpenStudy (anonymous):

I need 16/84 reduced I know how to do it but i couldn't figure it out Could you explain and answer I need 16/84 reduced I know how to do it but i couldn't figure it out Could you explain and answer @Mathematics

OpenStudy (anonymous):

You find a number that will go into both 16 and 84.. since they are even you can always begin with 2. 16 / 2 8 -- = -- 84 / 2 42 but 2 goes into both of those 8 /2 4 -- = -- 42 /2 21 Or if you knew 4 went into both at the beginning you could have 16 / 4 4 -- = -- 84 / 4 21

OpenStudy (anonymous):

Find a common factor of both the numerator and denominator, and divide them by it. For example both 16 and 84 can be divided by 4 to become 4/21.

OpenStudy (anonymous):

thanks so much